OpenStudy (linn99123):

A circular pan has a diameter of 8 inches. What is the exact area of the pan? A. 64π square inches B. 8π square inches C. 16π square inches D. 256π square inches

OpenStudy (fibonaccichick666):

getting people to supply you answers is not learning.

OpenStudy (akashdeepdeb):

Area is given by the equation pi*r*r aND r = DIAMETER/2 You have been given the diameter. Calculate the area! :D

OpenStudy (linn99123):

so its 8?

OpenStudy (akashdeepdeb):

Not quite. Diameter = 2* radius  So you have been given the diameter = 8 inches Thus the radius has to be = 4 inches Area of circle is given by = pi * radius * radius pi \(\approx\) 3.14 [Universal value of pi] radius = 4 inches So area = 3.14 * 4 * 4 = 16 * 3.14 = 16*pi = 16\(\pi\) [Getting the answer is not important, getting the concept is] [Go over through the steps again, to do these kinds of questions yourself :D ] Understood? :)